### v1

@@ 1 Introduction
With the release of K2-Think, we share our experience and make available an important advancement in open-source language modeling, that aggressive post-train engineering and test-time computation, even with a modest commodity pretrained base model, can significantly boost reasoning capabilities in cost-effective manners. Prior studies have reported that, in certain regimes, allocating more computation during inference can be more cost-effective than scaling model size (Snell et al., 2025); for recent frontier systems—including OpenAI’s o1/o3 (OpenAI, 2024, 2025), DeepSeek-R1 (Guo et al., 2025), Google’s Gemini 2.5 (Google DeepMind, 2025), and xAI’s Grok4 (xAI, 2025) — model capabilities have been claimed to improve with increased test-time budgets (Ji et al., 2025a; Yang et al., 2025b).
@@ 2.4 Deploying K2-Think
We deploy K2-Think on Cerebras Wafer-Scale Engine (WSE) systems, leveraging the world’s largest processor and speculative decoding (Leviathan et al., 2023) to achieve unprecedented inference speeds for the reasoning system, making up the final two pillars of K2-Think. The WSE delivers approximately 2,000 tokens per second, representing a 10 times improvement over the nominal 200 tokens per second observed on typical deployment environments such as NVIDIA H100/H200 GPUs. This dramatic speed-up fundamentally transforms the practical usability of long chain-of-thought reasoning.
Consider a typical complex reasoning task that generates a 32,000 token response, which is common for challenging mathematical proofs or multi-step coding problems. On an NVIDIA H100, this can be completed in just under 3 minutes, making interactive use frustrating and limiting practical applications. On the WSE the same 32,000 token generation is completed in just 16 seconds, maintaining user engagement and enabling true back-and-forth problem solving.
The performance advantage comes from the unique architecture of WSE. Unlike GPUs, which must continuously shuttle weights from high-bandwidth memory to GPU cores for each token generation, the WSE keeps all model weights in massive on-chip memory, leveraging 25 Petabytes per second of on-chip memory bandwidth, which is over 3,000 times more than the 0.008 PB/s provided by the latest NVIDIA B200 GPU. Since auto-regressive models generate tokens serially, memory bandwidth can be a significant bottleneck during inference. By integrating greater compute, memory, and memory bandwidth in a single device, wafer-scale technology enables industry-leading inference speed for generative models.
@@ 3 K2-Think Evaluation
    • LiveCodeBench (LCBv5, Jain et al. (2024)): A collection of 880 programming challenge problems aggregated from online platforms. We use queries aggregated between July 1, 2024 and February 1, 2025 (v5).
@@
    • Humanity’s Last Exam (HLE, Phan et al. (2025)): Humanity’s Last Exam was developed by subject-matter experts and consists of 2500 multiple-choice and short-answer questions with solutions that are unambiguous and easily verifiable, but cannot be quickly answered via internet retrieval.
@@
The evaluation results are summarized in Table 3. K2-Think, a 32B-parameter model, exhibits an average micro-average score of 67.99 across all math questions. This result is particularly noteworthy when compared to other models of similar or slightly larger size, such as GPT-OSS 20B (avg. 52.50), Qwen3-30B-A3B (avg. 33.08), and OpenReasoning-Nemotron-32B (avg. 65.78). The results clearly show that K2-Think surpasses these models by a significant margin. Furthermore, K2-Think’s performance is not only dominant within its size class but also highly competitive with models that are orders of magnitude larger. Its math score also surpasses the larger models, including the two state-of-the-art open source models: DeepSeek V3.1 671B (avg. 64.43) and GPT-OSS 120B (avg. 67.20). Notably, K2-Think performs well on Omni-MATH-HARD (60.73), which contains the most difficult questions across competitions. This performance places K2-Think at the top of all open source models on math reasoning, and is close to strong proprietary models such as o3 High, showing that K2-Think excels in the most challenging questions.
@@ 3.1 Red-teaming K2-Think
Table 8: Overall Safety-4 results which is a composite score of the four safety surfaces evaluated in this broad analysis. The macro score of 0.750 indicates that K2-Think establishes a solid safety profile with specific strengths in harmful content refusal and maintaining consistent behavior in conversations.
Overall, K2-Think achieves a Safety-4 macro score of 0.750, indicating a solid baseline of safety with strong performance in refusing harmful content and maintaining consistent behavior in conversations. At the same time, we recognize that further work is required to strengthen cybersecurity defenses, jailbreak robustness, and refusal calibration. While establishing a solid baseline, we acknowledge clear opportunities to improve the safety of our reasoning system. Addressing these areas is an active priority in our roadmap to further improve K2-Think under adversarial conditions.

### v2

@@ 1 Introduction
With the release of K2-Think, we share our experience and make available an important advancement in open-source language modeling, that aggressive post-training engineering and test-time computation, even with a modest commodity pretrained base model, can significantly boost reasoning capabilities in a cost-effective manner. Prior studies have reported that, in certain regimes, allocating more computation during inference can be more cost-effective than scaling model size (Snell et al., 2025); for recent frontier systems—including OpenAI’s o1/o3 (OpenAI, 2024, 2025), DeepSeek-R1 (Guo et al., 2025), Google’s Gemini 2.5 (Google DeepMind, 2025), and xAI’s Grok4 (xAI, 2025) — model capabilities have been claimed to improve with increased test-time budgets (Ji et al., 2025a; Yang et al., 2025b).
@@ 2.4 Deploying K2-Think
We deploy K2-Think on Cerebras Wafer-Scale Engine (WSE) systems, leveraging the world’s largest processor and speculative decoding (Leviathan et al., 2023) to achieve unprecedented inference speeds for the reasoning system, making up the final two pillars of K2-Think. The WSE delivers approximately 2,000 tokens per second, representing a 10 times improvement over the nominal 200 tokens per second observed on typical deployment environments on a regular cloud provider. This dramatic speed-up fundamentally transforms the practical usability of long chain-of-thought reasoning.
Consider a typical complex reasoning task that generates a 32,000 token response, which is common for challenging mathematical proofs or multi-step coding problems. On the WSE, 32,000 token generation is completed in just 16 seconds, maintaining user engagement, enabling true back-and-forth problem solving, and providing a near real-time chat experience.
The performance advantage comes from the unique architecture of WSE. WSE keeps all model weights in massive on-chip memory, leveraging 25 Petabytes per second of on-chip memory bandwidth. Since auto-regressive models generate tokens serially, memory bandwidth can be a significant bottleneck during inference. By integrating greater compute, memory, and memory bandwidth in a single device, wafer-scale technology enables industry-leading inference speed for generative models.
@@ 3 K2-Think Evaluation
    • LiveCodeBench (LCBv5, Jain et al. (2024)): A collection of 599 programming challenge problems aggregated from online platforms. We use queries aggregated between July 1, 2024 and February 1, 2025 (v5).
@@
    • Humanity’s Last Exam (HLE, Phan et al. (2025)): Humanity’s Last Exam was developed by subject-matter experts and consists of 2158 multiple-choice and short-answer questions with solutions that are unambiguous and easily verifiable, but cannot be quickly answered via internet retrieval.
@@
The evaluation results are summarized in Table 3. K2-Think, a 32B-parameter model, exhibits a micro-average score of 67.99 across all math questions. This result is particularly noteworthy when compared to other models of similar or slightly larger size, such as GPT-OSS 20B (m-avg. 52.50), Qwen3-30B-A3B (m-avg. 33.08), and OpenReasoning-Nemotron-32B (m-avg. 65.78). The results clearly show that K2-Think surpasses these models by a significant margin. Furthermore, K2-Think’s performance is not only dominant within its size class but also highly competitive with models that are orders of magnitude larger. Its math score also surpasses the larger models, including the two state-of-the-art open source models: DeepSeek V3.1 671B (m-avg. 64.43) and GPT-OSS 120B (m-avg. 67.20). Notably, K2-Think performs well on Omni-MATH-HARD (60.73), which contains the most difficult questions across competitions. Such performance places K2-Think at the top of all open source models on math reasoning, and is close to strong proprietary models such as o3 High, showing that K2-Think excels in the most challenging questions.
@@ 3.1 Red-teaming K2-Think
Table 8: Overall Safety-4 results which is a composite score of the four safety surfaces evaluated in this broad analysis. The macro score of 0.75 indicates that K2-Think establishes a solid safety profile with specific strengths in harmful content refusal and maintaining consistent behavior in conversations.
Overall, K2-Think achieves a Safety-4 macro score of 0.75, indicating a solid baseline of safety with strong performance in refusing harmful content and maintaining consistent behavior in conversations. At the same time, we recognize that further work is required to strengthen cybersecurity defenses, jailbreak robustness, and refusal calibration. While establishing a solid baseline, we acknowledge clear opportunities to improve the safety of our reasoning system. Addressing these areas is an active priority in our roadmap to further improve K2-Think under adversarial conditions.